PLEASE HELP!!
5^t-1 • 3^2t = 4^t+3

Answer by Alan3354(69443)

5^t-1 • 3^2t = 4^t+3
===================
Use parentheses to clarify. They're free.
If you mean:
5^(t-1) • 3^(2t) = 4^(t+3)
(t-1)*log(5) + 2t*log(3) = (t+3)log(4)
==================
Also, say what you want to do.
Probably solve for t.